40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Dec 4 10:51:10 2012, Steve, Update, SUS, PRM damping restored 
    Reply  Mon Dec 17 14:13:55 2012, Jenne, Update, SUS, PRM oplev gains restored 
Message ID: 7838     Entry time: Mon Dec 17 14:13:55 2012     In reply to: 7781
Author: Jenne 
Type: Update 
Category: SUS 
Subject: PRM oplev gains restored 

Quote:

PRM oplev gains set to zero from PIT 0.15 and YAW -0.3 and damping restored

 Put them back to normal.

ELOG V3.1.3-